Differential Analysisfor a Discontinued Product

A condensed income statement by product line for Healthy Beverage Inc. indicated the following for Fruit Cola for the past year:

Sales $234,600
Cost of goods sold 112,000
Gross profit $122,600
Operating expenses 145,000
Loss from operations $(22,400)

It is estimated that 12% of the cost of goods sold represents fixed factory overhead costs and that 23% of the operating expenses are fixed. Because Fruit Cola is only one of many products, the fixed costs will not be materially affected if the product is discontinued.

a. Prepare a differential analysis dated January 5 to determine whether Fruit Cola should be continued (Alternative 1) or discontinued (Alternative 2). If an amount is zero, enter "0". Use a minus sign to indicate a loss.

Differential Analysis
Continue Fruit Cola (Alt. 1) or Discontinue Fruit Cola (Alt. 2)
January 5
Continue Fruit Cola (Alternative 1) Discontinue Fruit Cola (Alternative 2) Differential Effect on Income (Alternative 2)
Revenues $ $ $
Costs:
Variable cost of goods sold
Variable operating expenses
Fixed costs
Income (Loss) $ $ $
Answer #1

Answer

Differential Analysis
Continue Fruit Cola (Alt. 1) or Discontinue Fruit Cola (Alt. 2)
Jan-05
Continue Fruit Cola (Alternative 1) Discontinue Fruit Cola (Alternative 2) Differential Effect on Income (Alternative 2)
Revenues $                                             234,600 $                                                            -   $                                                             (234,600)
Costs:
Variable cost of goods sold $                                               98,560 $                                                            -   $                                                               (98,560)
Variable operating expenses $                                             111,650 $                                                            -   $                                                             (111,650)
Fixed costs $                                               46,790 $                                                    46,790 $                                                                         -  
Income (Loss) $                                             (22,400) $                                                  (46,790) $                                                               (24,390)

Working Notes:

Variable cost of goods sold = $112,000 * 0.88 = 98,560

Variable operating expenses = $145,000 * 0.77 = 111,650

Fixed costs = ($112,000 * 0.12)+($145,000 * 0.23) = 46,790

In case of any doubt, please comment.